taking this math class online and having tons of problems. thanks for your help
remove and factor by 1
3w-27/3w
-33+11w/5 / 9w-27/25
solve for y
y-6/y+6=4/7
3- a-17/a+5=a^2-3/a+5

Answer by edjones(8007)

(y-6)/(y+6)=4/7 I assume that this is the problem.
7(y-6)=4(y+6) Multiply each side by the LCM 7(y+6) to eliminate the fractions.
7y-42=4y+24
3y=66
y=22
.
All the other problems are similar to this. Just eliminate the fractions by multiplying by the LCM and solve them.
.
Ed